The neutralization reaction between hydrochloric acid (HCI) and calcium hydroxide [Ca(OH)2] produces calcium chloride )CaCl2) and water.


2hCl(aq) + Ca(OH)2(s) 🠖 2H2O(l) + CaCl2(aq)


If 0.75 grams of Ca(OH)2 is used in the reaction, determine the following:


a. Mass of HCl (MW=36.64 g/mol)needed to completely react with Ca(OH)2. The molar mass of Ca(OH)2 is 74.10 g/mol.


b. Mass of CaCl2 is (MW=110.98 g/mol) formed in the reaction.



1
Expert's answer
2022-04-28T07:36:03-0400

"n(Ca(OH)2)=0.75\/74.10=0.01 mol, m(HCl)=2*0.01*74.10=1.5 g; m(CaCl2)=110.98*0.01= 1.12 g"